CONTINUUM ABSORPTION IN THE REGION OF THE O_{2} VIBRATIONAL FUNDAMENTAL BAND IN O_{2} AND CO_{2} MIXTURES AT TEMPERATURES RANGING FROM 220 K TO 296 K

\underline{YU. A. BARANOV}, Institute of Experimental Meterology, Taifun Research and Production Association, Obninsk, 249020 Russia; A. A. VIGASIN, Obukhov Institute of Atmospheric Physics, Pyzhevsky per 3, Moscow, 109017 Russia; W. J. LAFFERTY, and G. T. FRASER, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD. 20899, USA.

Infrared spectra of O_{2} and CO_{2} mixtures were studied in the laboratory between 1100 cm^{−1} and 1800 cm^{−1} for temperatures from 220 K to 296 K and total pressures up to 500 kPa using a Fourier-transform infrared spectrometer and a 2-meter-long, multiple-pathlength, White cell operating at 84 m. The intensity of the O_{2} continuum is strikingly enhanced by the addition of CO_{2}. This result contrasts with our observations on the CO_{2} monomer and dimer lines, which show negligible enhancement upon the addition of O_{2}. The shape of the O_{2} fundamental band sharpens with the addition of CO_{2}, and at lower temperatures apparent P, Q, and R-branch features appear, which are attributed to a CO_{2}–O_{2} complex. The observed CO_{2} enhancement of the O_{2} continuum absorption in the mid-infrared is consistent with our similar observations on the 1.27 μm near-infrared, magnetic-dipole band of O_{2} ^{a}. Our observations agree with the theoretical calculations by Brown and Tipping ^{b} in which they predicted significant enhancements by H_{2}O vapor of the continuum absorption of N_{2} and O_{2} for the vibrational fundamentals, attributed to the large electric dipole moment of H_{2}O. We note that CO_{2}, like H_{2}O, has a large electrostatic moment, the electric quadrupole moment.

aG. T. Fraser and W. J. Lafferty, J. Geophys. 106 31749 (2001).

bA. Brown and R. H. Tipping, private communication.